# Dark-Matter-Deficient Galaxies from Collisions: A New Probe of Bursty Feedback and Dark Matter Physics

### v2

@@ 5 Discussion and Conclusion
Our controlled simulations show that energy injection weakens the progenitors’ baryonic binding, elevating DMDG formation through collisions. Repeated potential fluctuations, as expected from bursty stellar feedback, may reduce $|E_{\rm bind}|$ by only 15%, yet in two thirds of the simulations, the DMDG masses are enhanced by over 100%. In contrast, SIDM creates cores without lowering $|E_{\rm bind}|$ and has a negligible effect on DMDG masses.
@@
In the regime of high mass ratios and low collision velocities, close encounters between satellite and host galaxies have been proposed as a primary channel for producing DMDGs within host systems (J. Moreno et al., 2022; D. Yang et al., 2020; Y. Jing et al., 2019; G. Ogiya, 2018). In such interactions, however, tidal stripping acts on both the dark matter and baryonic components of the satellite. If baryonic feedback is too strong, it can make the baryons too diffuse and easily stripped, hindering the formation of DMDGs. In contrast, SIDM can generate cores without significantly expanding the stellar distribution, offering a more efficient mechanism for DMDG formation (D. Yang et al., 2020). The relative abundance of DMDGs in the field versus those in host halos thus provides a potential observational handle to distinguish between core formation driven by baryonic feedback and that induced by SIDM.
@@
The distinct formation pathways of DMDGs lead to contrasting observational signatures. Systems produced through tidal stripping resemble ordinary satellites but typically exhibit lower surface brightness. In contrast, those formed from post-collision tidal debris contain a substantial stellar component with nearly uniform ages and metallicities, created in a short burst at the time of the encounter. Another class of baryon-dominated systems is the tidal dwarf galaxy, which forms from tidal-tail gas ejected during a strong interaction involving a massive disk (F. Bournaud & P. A. Duc, 2006; G. Gentile et al., 2007; S. Kaviraj et al., 2012; J. Zaragoza-Cardiel et al., 2024). These galaxies are typically embedded within tidal tails and generally lack bright globular clusters (J. Lee et al., 2024).
@@
Appendix A Star formation in gas remnants
@@
$b_{T}\equiv\frac{\Sigma_{\rm gas}}{\sigma_{v,\rm gas}^{2}+c_{s}^{2}}\propto\left(\frac{\tau_{\rm dyn}}{\tau_{\rm ff}}\right)^{2},$ (A1)
@@
Figure 6: Post-collision star formation. Mass evolution of the most massive DMDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened binding case ($\gamma=0.1$), efficient star formation yields a more massive DMDG after the collision, with continued activity beyond 2 Gyr steadily depleting the gas reservoir. For $\gamma=1$, star formation is much less efficient, leaving a substantial gas fraction even at 5 Gyr. In both scenarios, the total DMDG mass increases only modestly, from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) case.
@@
Appendix B Stability of SIDM cores
Figure 7: Stability test of SIDM halos. The progenitor system in the 1c simulation is first evolved under SIDM ($\sigma/m=20~\rm cm^{2}~g^{-1}$) for 2 Gyr (dotted lines), when a dark matter core rapidly forms and reaches a stable configuration. SIDM interactions are then turned off and the system is further evolved for 2 Gyr (solid lines). Both the dark matter (red) and stellar (blue) components remain stable, confirming that the SIDM-induced core is long-lived and justifying the neglect of SIDM in the subsequent collisional simulations.
The SIDM benchmark in our study is implemented by incorporating cored halos obtained from a pre-simulation. Specifically, we use Gadget-2 to evolve an initial system consisting of a gas disk and dark matter under gravity and SIDM ($\sigma/m=20~\rm cm^{2}~g^{-1}$) for 2 Gyr. The final snapshot is then passed to Gadget-4 to initialize the hydrodynamical collisional simulation. This strategy is justified by the following considerations.
Figure 7 illustrates this verification for the progenitor system in the 1c simulation. During the first 2 Gyr, the halo evolves under SIDM: the dark matter core develops rapidly within the first $\sim$1 Gyr and then approaches a stable configuration. After 2 Gyr, SIDM is switched off and the simulation continues. Both the dark matter (red) and stellar (blue) components remain stable for at least another 2 Gyr. This demonstrates that neglecting SIDM in the collisional simulation is a valid approximation.
Appendix C Tabulated full simulation results
We adopt the gas cooling and star formation module provided in the Gadget-4 code. Radiative cooling is implemented through tabulated rates for primordial gas, while stars form stochastically from dense gas following the multiphase ISM prescription of V. Springel & L. Hernquist (2003). Although our study does not focus on the detailed dynamical processes within UDGs, the efficiency of gas cooling plays a key role in regulating their stability: a lower cooling rate is generally expected to promote stability. To demonstrate that our conclusions are robust against variations in the cooling rate, we present the full simulation results in Table 2 for the default cooling rate and in Table 3 for a case with the cooling rate reduced by half. In these tables, we identify DMDGs as systems with $M_{*}(r<10{\rm kpc})>10^{6}M_{\odot}$, reporting their abundance as well as the mass content of the most massive DMDGs in each simulation.
Figure 8 summarizes the reduced cooling case, in the same format as Figure 4 in the main text. All key features, most notably the enhancement in DMDG mass and star formation efficiency, remain unaffected. A comparison of the numerical results across the two tables shows that reduced cooling produces more DMDGs with lower characteristic masses. This indicates a potential degeneracy between the effects of gas cooling and weakened gravitational binding. The degeneracy can be resolved by the observed abundance of UDGs, which provides a means to constrain gas cooling (Y. Rong et al., 2017; R. F. J. van der Burg et al., 2017; A. Di Cintio et al., 2017; M. G. Jones et al., 2018).
@@ Appendix C Tabulated full simulation results
Table 2: Galaxy collision simulations with Gadget-4: initial collision configurations and properties of the resulting dark-matter-dominated galaxies (DMDGs).
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $W/E_{\rm bind}$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)
1a 1.5 0.15 $400$ 22 $14$ 1 8.75 1 0.38 1.49
1b 0.1 7.58 1 8.61 4.85
1c 1.5 0.15 400 2 14 SIDM 9.47 1 0.46 2.97
2a 2 0.2 $400$ 22 $14$ 1 14.73 4 0.10 2.13
2b 0.1 12.49 1 6.09 6.72
3a 2 0.2 $280$ 22 $14$ 1 14.73 2 4.89 3.70
3b 0.1 12.49 1 16.80 4.97
4a 1.0 0.1 $400$ $1.5$ 44 1 3.53 1 0.15 4.13
4b 0.1 2.87 1 0.30 5.54
5a 1.5 0.15 $400$ 22 44 1 7.17 1 6.96 8.32
5b 0.1 5.82 1 9.61 5.05
6a 1.5 0.22 $400$ 22 44 1 10.99 2 16.84 6.02
6b 0.1 9.28 1 20.70 7.03
7a 1.5 0.22 $450$ 22 44 1 10.99 2 2.39 7.74
7b 0.1 9.28 2 3.23 3.96
8a 1.5 0.15 $400$ 22 77 1 7.87 3 1.62 3.38
8b 0.1 6.47 1 11.15 3.91
9a 1.5 0.22 $400$ 22 77 1 11.93 3 4.66 4.42
9b 0.1 10.21 2 18.33 5.57
10a 1.5 0.22 $280$ 22 77 1 11.93 2 12.71 3.15
10b 0.1 10.21 1 25.90 7.27
11a 1.5 0.15 $280$ 22 $14$ 1 8.75 1 0.22 2.14
11b 0.1 7.58 1 8.38 5.20
12a 1.5 0.22 $400$ 22 $14$ 1 13.56 3 3.12 5.88
12b 0.1 11.68 1 13.15 5.23
13a 2 0.2 $400$ 22 44 1 11.72 1 14.64 6.60
13b 0.1 9.75 1 19.49 5.88
14a 2 0.2 $400$ 22 77 1 12.81 2 0.78 3.01
14b 0.1 10.83 1 13.97 5.53
15a 5 0.5 $500$ $3.5$ $14$ 1 78.10 3 13.81 4.98
15b 0.1 66.40 1 26.08 5.11
@@
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $W/E_{\rm bind}$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)

### v3

@@ VI Discussion and Conclusion
Our controlled simulations show that energy injection weakens the progenitors’ baryonic binding, elevating DMDG formation through collisions. Repeated potential fluctuations, as expected from bursty stellar feedback, may reduce $|E_{\rm bind}|$ by only 15%, yet in two thirds of the simulations, the DMDG masses are enhanced by over 100%. In contrast, SIDM creates cores without lowering $|E_{\rm bind}|$, resulting in heavier and lighter DMDGs from the $\gamma=1$ cases with similar chances.
@@
In the regime of high mass ratios and low collision velocities, close encounters between satellite and host galaxies have been proposed as a primary channel for producing DMDGs within host systems (J. Moreno et al., 2022; D. Yang et al., 2020; Y. Jing et al., 2019; G. Ogiya, 2018). We simulate the three scenarios under the same orbit and host halo, showing that both baryon feedback and SIDM can foster DMDG formation through enhanced tidal stripping, but with different capabilities. The relative abundance of DMDGs in the field versus those in host halos thus provides a potential observational handle to distinguish between core formation driven by baryonic feedback and that induced by SIDM.
@@
The distinct formation pathways of DMDGs may lead to contrasting observational signatures. Systems produced through tidal stripping resemble ordinary satellites but typically exhibit lower surface brightness. In contrast, those formed from post-collision tidal debris contain a substantial stellar component with nearly uniform ages and metallicities, created in a short burst at the time of the encounter. Another class of baryon-dominated systems is the tidal dwarf galaxy, which forms from tidal-tail gas ejected during a strong interaction involving a massive disk (F. Bournaud & P. A. Duc, 2006; G. Gentile et al., 2007; S. Kaviraj et al., 2012; J. Zaragoza-Cardiel et al., 2024). These galaxies are typically embedded within tidal tails and generally lack bright globular clusters (J. Lee et al., 2024).
@@
Appendix A Stability of SIDM cores
Figure 6: Left: The stability test of SIDM halos. The progenitor system in the 1c simulation is first evolved under SIDM ($\sigma/m=20~\rm cm^{2}~g^{-1}$) for 2 Gyr (dotted lines), when a dark matter core rapidly forms and reaches a stable configuration. SIDM interactions are then turned off and the system is further evolved for 2 Gyr (solid lines). Both the dark matter (red) and stellar (blue) components remain stable, confirming that the SIDM-induced core is long-lived and justifying the neglect of SIDM in the subsequent collisional simulations. Right: The tidal tripping effect for the SIDM core phase with and without SIDM effect. The orange lines are consistent with those in Figure 5. The green lines presents the same initial condition to the orange lines, but do not include the SIDM effect during the simulation process.
The SIDM benchmark in our study is implemented by incorporating cored halos obtained from a pre-simulation. Specifically, we use Gadget-2 to evolve an initial system consisting of a gas disk and dark matter under gravity and SIDM ($\sigma/m=20~\rm cm^{2}~g^{-1}$) for 2 Gyr. The SIDM simulation is based on a well-tested module developed in prior works (D. Yang & H.-B. Yu, 2022; D. Yang et al., 2023). Since Gadget-2 does not include radiative cooling and star formation modules, we pass the final snapshot from the pre-simulation to Gadget-4 for the hydrodynamical collision simulation. For computational efficiency, we adopt a relatively large SIDM interaction $\sigma/m=20~\rm cm^{2}~g^{-1}$, which leads to significant core formation within $\sim$ 2 Gyr. The resulting cored halo distribution can be interpreted as arising from lower cross sections with a longer evolution time ($t_{\rm evo}$), based on a degeneracy in which $t_{\rm evo}\sigma/m$ is roughly constant. For example, one can equivalently consider $\sigma/m=5~\rm cm^{2}~g^{-1}$ with an 8 Gyr pre-evolution.
In our high-velocity dwarf collision simulations, the effect of SIDM primarily goes into the initial conditions. This is because the time scale of the dwarf collisions is very short. In the 2a benchmark, for example, the two halos pass each other and separate by about 100 kpc in just 0.4 Gyr. At this point, the region where DMDGs form has become baryon-dominated, and SIDM does not play a role. Aside from this, we provide additional justifications for our approach below.
The left panel in Figure 6 illustrates this verification for the progenitor system in the 1c simulation. During the first 2 Gyr, the halo evolves under SIDM: the dark matter core develops rapidly within the first $\sim$1 Gyr and then approaches a stable configuration. After 2 Gyr, SIDM is switched off and the simulation continues. Both the dark matter (red) and stellar (blue) components remain stable for at least another 2 Gyr. These demonstrate that neglecting SIDM in the collisional simulation is a valid approximation.
In our low-velocity collision simulations, where the satellite evolves in the host potential for over 10 Gyr, we have consistently simulated the SIDM effect. Here, we compare simulation results with and without SIDM enabled during evolution. Figure 6 shows the bound dark matter (solid) and stellar (dashed) mass evolution with (orange) and without (green) SIDM. As expected, their evolution trajectories are very close to each other. SIDM with a cross section per mass $\sigma/m=5~\rm cm^{2}/g$ only slightly boosts the tidal stripping, demonstrating that the SIDM effect primarily goes into the initial condition, even for these cored systems in the host.
Appendix B Tidal effect
Figure 7: Additional contributing aspects of post collision gas condensation. The upper panels plot the ratio of the gas sound speed ($c_{s}$) to its velocity dispersion ($\sigma_{v}$). We analyze the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) benchmarks from Table 1 and present results at t=0.4t=0.4~Gyr, when DMDGs are about to form following the collision. In the region where gas collapse, we find $\sigma_{v}\gg c_{s}$, which suggests gravity dominates over internal pressure in driving the dynamics of the gas. The bottom panels illustrate an approximation of the tidal to self-gravitational acceleration ratio, defined as $a_{\rm tidal}/a_{\rm self}\sim\frac{\int\lambda_{\rm max}{\rm d}z}{\int G\Sigma{\rm d}z}$. The results are presented for the same benchmarks as in the upper panels. A bulk of the region shows magnitudes greater than 1, implying that a stronger tidal effect shapes the gas collapse.
The main text examined the impact of tidal forces on post-collision gas condensation. Here, we consider two additional factors, i.e., thermal pressure and self-gravity, and demonstrate that their effects are secondary at intermediate scales, where gas accumulates in regions likely to form DMDGs.
First, we show that thermal pressure ($\rho c_{s}^{2}$) is less significant than kinetic pressure ($\rho\sigma^{2}$). The upper panels in Figure 7 display maps of the sound speed to velocity dispersion ratio, $c_{s}/\sigma_{v}$, for the 2a and 2b benchmarks at t=0.4t=0.4~Gyr. The results are consistently below one in regions of gas collapse, confirming that kinetic pressure dominates over thermal pressure.
To show that the tidal winds effectively reshape the gas collapse, we plot an approximate quatity for the ratio of tidal to self-gravitational acceleration, defined as $a_{\rm tidal}/a_{\rm self}\sim\frac{\int\lambda_{\rm max}{\rm d}z}{\int G\Sigma{\rm d}z}$, where $\lambda_{\rm max}$ is the largest eigenvalue of the tidal tensor. The lower panels of Figure 7 present the results considering the same benchmarks and snapshot as in the upper panels. The results show $|a_{\rm tidal}/a_{\rm self}|>1$ in most areas, suggesting that there is a phase of tidal dominance before DMDGs form under the baryon’s self-gravity.
Appendix C Star formation in gas remnants
@@
$b_{T}\equiv\frac{\Sigma_{\rm gas}}{\sigma_{v,\rm gas}^{2}+c_{s}^{2}}\propto\left(\frac{\tau_{\rm dyn}}{\tau_{\rm ff}}\right)^{2},$ (C1)
@@
Figure 9: Mass evolution of the most massive DMDGs in the 2a ($\gamma=1$) and 2b ($\gamma=0.1$) simulations. In the weakened binding case ($\gamma=0.1$), efficient star formation yields a more massive DMDG after the collision, with continued activity beyond 2 Gyr steadily depleting the gas reservoir. For $\gamma=1$, star formation is much less efficient, leaving a substantial gas fraction even at 5 Gyr. In both scenarios, the total DMDG mass increases only modestly, from $11\times 10^{8}~\rm M_{\odot}$ ($2.3\times 10^{8}~\rm M_{\odot}$) to $15\times 10^{8}~\rm M_{\odot}$ ($2.6\times 10^{8}~\rm M_{\odot}$) in the $\gamma=0.1$ ($\gamma=1$) case.
@@
Appendix D Tabulated full simulation results
We adopt the gas cooling and star formation module provided in the Gadget-4 code. Radiative cooling is implemented through tabulated rates for primordial gas, while stars form stochastically from dense gas following the multiphase ISM prescription of V. Springel & L. Hernquist (2003). Although our study does not focus on the detailed dynamical processes within UDGs, the efficiency of gas cooling plays a key role in regulating their stability: a lower cooling rate is generally expected to promote stability. To demonstrate that our conclusions are robust against variations in the cooling rate, we present the full simulation results in Table 2 for a case with the cooling rate reduced by half. In these tables, we identify DMDGs as systems with $M_{*}(r<10{\rm kpc})>10^{6}M_{\odot}$, reporting their abundance as well as the mass content of the most massive DMDGs in each simulation.
Figure 10 summarizes the reduced cooling case, in the same format as Figure 4 in the main text. All key features, most notably the enhancement in DMDG mass and star formation efficiency, remain unaffected. A comparison of the numerical results across the two tables shows that reduced cooling produces more DMDGs with lower characteristic masses. This indicates a potential degeneracy between the effects of gas cooling and weakened gravitational binding. The observed abundance of UDGs may help to distinguish such differences, providing a means to constrain gas cooling (Y. Rong et al., 2017; R. F. J. van der Burg et al., 2017; A. Di Cintio et al., 2017; M. G. Jones et al., 2018).
@@ Appendix D Tabulated full simulation results
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $|E_{\rm bind}|$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) $E_{0}$ ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)